OpenStudy (anonymous):

Find the solution set of 4^2 + 7x - 2 So far I have (4x^2 - 1x) (8x - 2) = 0 This is where I'm stuck. I have to pull 1x from the first set of parenthesis?

OpenStudy (anonymous):

Or -1x? Not sure what to do next.

OpenStudy (anonymous):

x(4x-1) 2(4x-1) (4x-1) (x+2) 4x-1=0 4x=1 x=1/4 x+2=0 x=-2

OpenStudy (anonymous):

{1/4, -2}